OpenStudy (anonymous):

The probability that Emma has heart disease is 0.6. She decides to take a test for the disease that is 80% accurate. What is the probability that Emma has heart disease and the test predicts heart disease?

OpenStudy (anonymous):

i need an explanation too

OpenStudy (dumbcow):

Ok here we have 2 independent events, you can find prob that both are true by multiplying the probabilities \[P(A and B) = P(A)*P(B)\] \[= P(heart disease) * P(test accurate)\] \[= 0.6* 0.8\]